# Modeling the formation of N2 and CH4 frost on Pluto’s slopes

### v1

@@ Abstract
Context. The equatorial region of Cthulhu, revealed by New Horizons, appears generally dark and largely devoid of volatiles due to its low-albedo surface. However, localized bright patches, interpreted as CH4 frost, are observed on crater rims and slopes.
Aims. Previous studies suggested that these frosts could result from the peculiar insolation driven by the geometry of these slopes, but this has never been quantitatively tested. We aim to investigate the origin, stability, and potential role in Pluto’s volatile cycle of these localized frost deposits.
Methods. We implemented a new sub-grid-scale slope parameterization in the Pluto Volatile Transport Model, which accounts for the specific solar irradiation and resulting surface and subsurface temperatures on sloped terrains. This parameterization also allows the condensation and sublimation of volatiles (either N2 or CH4) on slopes, including the effect of large-scale transport of these species, which is key to determining the amount of frost that forms or disappears.
Results. Our simulations reproduce the observed CH4 frost on North-facing slopes as seasonal deposits currently sublimating, predict perennial CH4 frost on South-facing slopes, and show that slope microclimates should not alter global volatile cycles.
Conclusions. Seasonal and perennial N2 and CH4 frosts can form across Pluto’s slopes, even in its darkest and warmest regions, due to the locally reduced sunlight received on inclined terrain. Nevertheless, despite Pluto’s abundance of sloped surfaces, slope microclimates appear to have only a minor effect on the planet’s global volatile cycles.
@@ 1 Introduction
The flyby of Pluto and its satellites by the NASA’s New Horizons spacecraft in July 2015 revealed a geologically and climatically dynamic world, with diverse surface ices and a tenuous but complex atmosphere (Stern et al., 2015; Grundy et al., 2016; Gladstone et al., 2016; Moore et al., 2016). The Multispectral Visible Imaging Camera (MVIC) and the Linear Etalon Imaging Spectral Array (LEISA) composing the Ralph spectrograph onboard the spacecraft detected a spatially heterogeneous distribution of volatiles mainly composed of N2, CH4, and CO onto the water ice bedrock (Protopapa et al., 2017; Schmitt et al., 2017; Earle et al., 2018b) that we summarized here. The most prominent N2 ice deposit on Pluto’s surface is the kilometer-thick sheet- mixed with CH4 and CO ice- in Sputnik Planitia (Stern et al., 2015; Grundy et al., 2016). The rest of the planet exhibits a clear latitudinal stratification between CH4-rich and N2-rich terrains. The North Pole is covered by relatively pure and bright CH4-rich ice. The plains between 55°N–70°N and 25°N–35°N are covered in mixtures of CH4 ice - possibly including CO and N2 ice- with N2 ice deposits filling the bottom of some craters and depressions. In between (35°N–55°N), the composition of the surface is dominated by N2-rich ice deposits. The equatorial regions of Pluto ($\pm$25° latitude) exhibits a strong longitudinal variability in the apparence and composition of the surface (Buratti et al., 2017; Schmitt et al., 2017; Moore et al., 2018), with the presence of the N2-ice rich deposit in Sputnik Planitia, the relatively pure CH4-rich ice deposits in the Bladed Terrain Deposit, and the dark (albedo of $\sim$ 0.1 Hofgartner et al., 2023), overall volatile free plains in the Cthulhu region (20°N- 20°S, 0°E- 180°E, also known as the Belton region) .
@@
These observations, combined with numerical models such as volatile transport models (e.g., Hansen & Paige, 1996; Young, 2012, 2013; Bertrand & Forget, 2016; Bertrand et al., 2018, 2019), or Global Climate Models (GCMs; e.g., Zalucha & Michaels, 2013; Toigo et al., 2015; Forget et al., 2017) show that the dynamics of these ice deposits are driven by latitude, elevation, seasonal evolution, and long-term orbital cycles (see a review in Moore, 2021; Young et al., 2021). For example, Bertrand & Forget (2016) and Bertrand et al. (2018) showed that the accumulation of N2 ice occurs within Sputnik Planitia as a result of the lower elevation in the basin, leading to a stronger thermal infrared cooling balanced by a stronger condensation temperature of N2, leading to a higher condensation rate than at any other location. Bertrand et al. (2019) showed that the orbital configurations of Pluto in the last Earth million years would have led to a net accumulation of CH4-rich deposits in the equatorial regions of Pluto, and seasonal accumulation elsewhere. The persistent absence of perennial ice in Cthulhu’s equatorial region may stem from its exceptionally low albedo compared to the rest of Pluto’s surface , which enhances solar heating of the high-inertia water-ice substrate. This stored heat is then released during colder seasons, keeping surface temperatures higher than elsewhere and thereby inhibiting CH4 frost accumulation (Earle et al., 2018a). In contrast, the high albedo of the seasonal CH4 deposits at mid- and high latitudes favors surface cooling in winter, promoting N2 condensation and contributing to the formation of N2-rich terrains at these latitudes (Earle et al., 2018a; Bertrand et al., 2019). Note that while these models successfully capture the first-order mechanisms controlling the large-scale ice distribution observed by New Horizons, accurately quantifying the amounts and rates of sublimation or condensation remains challenging because key surface and thermal parameters (e.g., CH4 ice albedo, water-ice thermal inertia) are only loosely constrained by the available observations (Young et al., 2021).
@@
Recent analyses using high-resolution measurements by Ralph/MVIC and the Long-Range Reconnaissance Imager (LORRI) instruments revealed the presence of bright N2 and CH4 ice-rich deposits on crater floors and rims and mountain tops (Bertrand et al., 2020; Earle et al., 2022). More specifically, Bertrand et al. (2020) highlighted the presence of bright CH4 frost on the North-facing crater rims and slopes in the Cthulhu region; and on the top of Pigafetta and Elcano Montes. Using high-resolution modeling, they showed that methane deposits found at the top of these mountains form from a circulation-induced enrichment of gaseous methane a few kilometers above Pluto’s plains that favors methane condensation at mountain summits. Yet, their model can not be used to study the formation of CH4 deposits on crater rims, given the too-thin resolution of these craters compared to the model’s resolution. They suggested that these deposits found on the craters’ flanks are actually linked to the specific microclimates of these slopes. Given Pluto’s tenuous atmosphere, sloped terrains receive significantly different amounts of solar insolation compared to flat surfaces during winter/spring (see their Figures S8, S9). This creates localized cold microclimates on the slope that can favor the accumulation of volatiles and mitigate their sublimation. Hence, they suggested that the frosts observed on Cthulhu crater’s rims are seasonal frost formed during the previous winter and should now be subliming. Earle et al. (2022) extended this work to other regions of Pluto and found a latitudinal dependence in the orientation of CH4 frost on crater rims. They showed that methane frost is detected on North-facing slopes in the equatorial regions of Pluto ($\pm$ 15° latitude), and equatorward-facing slopes at higher latitudes (between 15° N and 38°N), and then on all of the crater rims at higher latitudes. Although they share the same hypothesis as Bertrand et al. (2020) to explain the presence of methane frost at low latitudes on North-facing slopes, they note that the change in preferential orientation in regions around 18°N could be explained by different formation timescales (e.g., seasonal at low latitudes, perennial at high latitudes). Finally, for N2 frosts, they showed that they are preferentially observed on the crater’s floor up to 38°latitude, and on equatorward-facing slopes at higher latitudes.
@@
So far, no numerical model has been able to address the frost formation hypotheses proposed by Bertrand et al. (2020) and Earle et al. (2022). Indeed, these frosts and their associated slope microclimates occur on spatial scales of only a few kilometers Bertrand et al. (2020), which is far smaller than the typical resolution of GCMs (e.g., hundreds of kilometers, Bertrand et al., 2019). Local models, such as 1D radiative equilibrium models, are also inadequate, as the stability of methane frost depends on its atmospheric mixing ratio, which is itself governed by large-scale meteorology that cannot be captured by these simplified models. This numerical gray zone has prevented further investigation into the dynamics of methane frosts on Pluto’s slopes so far.
@@
A similar problem was encountered by the Martian community with the modeling of water frost on tropical slopes. Due to its thin atmosphere transparent to visible radiation, slope-driven microclimates are significant on Mars. These slope microclimates allow the formation of CO2 frost (an analogue for N2 on Pluto) and water frost (an analogue for CH4 on Pluto) at subtropical latitudes, up to 20–30 degrees equatorward of the seasonal polar caps (e.g., Vincendon et al., 2010b; Lange et al., 2024). The same numerical limitations described above applied when attempting to model these low-latitude frosts on crater walls. Recently, Lange et al. (2023) introduced a sub-grid parameterization of slope microclimates into the Mars Planetary Climate Model. The principle of this parameterization is illustrated in Figure 2. For each grid cell of the global Mars climate model, the terrain is decomposed into a distribution of sloped surfaces (defined by characteristic slope angles) and a flat surface. On each sub-grid terrain, the surface energy balance and the condensation/sublimation of volatiles are computed, assuming all terrains interact with the same atmospheric column. The near-surface atmosphere then experiences an average effect of these sub-grid microclimates. This parameterization successfully reproduced the seasonal dynamics of CO2 and water frosts at low latitudes on Mars.
In this paper, we propose to adapt the slope microclimate parameterization developed by Lange et al. (2023) to the case of Pluto. Specifically, the objectives of this work are to:
  1. 1.
    Describe the seasonal evolution of methane and nitrogen frost on crater slopes over a Pluto year, with particular attention to the sensitivity to thermophysical and optical properties of the ices;
  2. 2.
    Discuss the nature of the frosts observed on crater flanks and slopes by New Horizons;
  3. 3.
    Discuss the potential impact of slope microclimates on Pluto’s global climate.
@@
To this end, we describe the parameterization in section 2. The dynamics of methane and nitrogen frost on sloped terrains, and the nature of the frosts observed by New Horizons are presented in section 3. The potential global climatic implications of slope microclimates are discussed in section 4.1. We draw our conclusions in section 5.
@@ 2.1 Frost identification and statistics
In order to compare our model with the observations of frost on slopes, we first need to determine the typical slope angle and azimuth at which frost is observed. While the mapping of N2/CH4 frosts has already been made in previous studies (e.g., Schmitt et al., 2017; Earle et al., 2018b), we focus here on the Cthulhu region as it is an unique location where frost is only observed on slopes and not on the dark plains of the regions (Bertrand et al., 2020), suggesting the crucial role of slope microclimates -and thereby slope angle and azimuth- in the formation of these frosts. As in Bertrand et al. (2020), we first map the location of bright deposits on LORRI observations (Hofgartner et al., 2023). As some of the bright locations might be linked to direct illumination, we then compare the locations of these bright deposits to the CH4 band depth map and phase index from the LEISA instrument (Schmitt et al., 2017; Gabasova et al., 2021) to derive the nature of these bright surface -frost or not, and if so, if it is pure N2, CH4 diluted in N2 ice or a CH4-rich ice phases-. Finally, we determine the slope angle and azimuth at which ice is observed using the Digital Evolution Model (DEM, with a 300 m resolution per pixel) from Schenk et al. (2018).
@@ 2.2 The Pluto Volatile Transport Model
For this study, we use the Pluto Volatile Transport Model (VTM) described in Bertrand & Forget (2016) (see their Methods section), which has been employed in previous investigations of ice dynamics on Pluto (Bertrand & Forget, 2016; Bertrand et al., 2018, 2019). In short, this model computes the surface radiative budget -assuming that the atmosphere is transparent at all wavelengths-, the exchange of heat with the subsurface by conduction, the exchange of heat with the surface by latent heat, and the exchanges of volatile (N2 and CH4) with the atmosphere. To represent horizontal transport after surface exchanges, the mixing ratios of atmospheric species are forced towards their global mean value, using a timescale of 107 s for CH4, 1 s for N2 estimated from 3D GCM studies (Bertrand & Forget, 2016; Forget et al., 2017).
Simulations are performed with a horizontal grid with a grid-point spacing of 1° by latitude and longitude ($\sim$ 21 km at the equator), and a vertical grid dividing the atmosphere into 27 levels ranging from a few meters close to the surface up to 140 km. The initial distribution of the N2 and CH4 reservoirs is obtained following the 30 million years of volatile ice evolution performed by Bertrand et al. (2018) and Bertrand et al. (2019): an infinite amount of N2 ice is placed within Sputnik Planitia, and an infinite reservoir of CH4 ice is placed in the Bladed Terrain Deposits (15°S-15°N, 140°W-15°E).
The emissivity and albedo of such ice deposits are tuning parameters of the model, and set to have a pressure and CH4 volume mixing ratio consistent with the measurements during the New Horizons fly-by (see section 2.4). The albedo of the frost/ice-free terrains are set to observed values from Hofgartner et al. (2023) and their emissivity to 1. The diurnal thermal inertia of the bare-surface, N2 ($\rm{TI}_{\rm{N}_{2},\rm{d}}$) and CH4 ($\rm{TI}_{\rm{CH}_{4},\rm{d}}$) are set by default to 9 J m-2 K-1 s-1/2 following low-thermal inertia derived by Lellouch et al. (2011) and Bertrand et al. (2025). The seasonal thermal inertia of the bedrock $\rm{TI}_{\rm{bedrock},\rm{s}}$ is another tuning parameter. The albedo of a surface is changed when ice/frost is present at the surface following the albedo-feedback parameterization of Bertrand et al. (2020).

### v2

# Modeling the formation of N2 and CH4 frost on the Pluto slopes

@@ Abstract
Context. The equatorial region of Cthulhu as revealed by New Horizons appears to be generally dark and largely devoid of volatiles because its surface albedo is low. Localized bright patches, however, which are interpreted as CH4 frost, are observed on crater rims and slopes.
Aims. Previous studies suggested that these frosts might result from the peculiar insolation driven by the geometry of these slopes, but this has never been tested quantitatively. We investigated the origin, stability, and potential role of these localized frost deposits in the volatile cycle of Pluto.
Methods. We implemented a new subgrid-scale slope parameterization in the volatile transport model for Pluto, which accounts for the specific solar irradiation and the resulting surface and subsurface temperatures on sloped terrains. This parameterization also allows the condensation and sublimation of volatiles (either N2 or CH4) on slopes, including the effect of large-scale transport of these species. This is key to determining the amount of frost that forms or disappears.
Results. Our simulations reproduce the observed CH4 frost on north-facing slopes as seasonal deposits that currently sublimate, predict perennial CH4 frost on south-facing slopes, and show that the slope microclimates are not expected to alter global volatile cycles.
Conclusions. Seasonal and perennial N2 and CH4 frosts can form on the Pluto slopes, even in its darkest and warmest regions, because the locally sunlight received on inclined terrain is reduced. Despite the abundance of sloped surfaces on Pluto, the slope microclimates still only appear to have a minor effect on the global volatile cycles of the planet.
@@ 1 Introduction
The flyby of Pluto and its satellites by the NASA New Horizons spacecraft in July 2015 revealed a geologically and climatically dynamic world, with diverse surface ices and a tenuous but complex atmosphere (Stern et al., 2015; Grundy et al., 2016; Gladstone et al., 2016; Moore et al., 2016). The Multispectral Visible Imaging Camera (MVIC) and the Linear Etalon Imaging Spectral Array (LEISA) composing the Ralph spectrograph on board the spacecraft detected a spatially heterogeneous distribution of volatiles that are mainly composed of N2, CH4, and CO onto the water-ice bedrock (Protopapa et al., 2017; Schmitt et al., 2017; Earle et al., 2018b) that we summarize here. The most prominent N2 ice deposit on the Pluto surface is the kilometer-thick sheet that is mixed with CH4 and CO ice in Sputnik Planitia (Stern et al., 2015; Grundy et al., 2016). The rest of the planet exhibits a clear latitudinal stratification between CH4-rich and N2-rich terrains. The north pole is covered by relatively pure and bright CH4-rich ice. The plains between 55°N–70°N and 25°N–35°N are covered in mixtures of CH4 ice (possibly including CO and N2 ice), with N2 ice deposits filling the bottom of some craters and depressions. Between 35°N–55°N, the composition of the surface is dominated by N2-rich ice deposits. The equatorial regions of Pluto (defined here as $\pm$15° latitude) exhibit a strong longitudinal variability in the appearance and composition of the surface (Buratti et al., 2017; Schmitt et al., 2017; Moore et al., 2018), with the N2-ice rich deposit in Sputnik Planitia, the relatively pure CH4-rich ice deposits in the bladed terrain deposit, and the dark (albedo of $\sim$ 0.1 Hofgartner et al., 2023) overall volatile-free plains in the Cthulhu region (defined here as the region between 20°N and 20°S, 0°E and 180°E, also known as the Belton region) .
@@
These observations, combined with numerical models such as volatile transport models (e.g., Hansen and Paige, 1996; Young, 2012, 2013; Bertrand and Forget, 2016; Bertrand et al., 2018, 2019), or global climate models (GCMs; e.g., Zalucha and Michaels, 2013; Toigo et al., 2015; Forget et al., 2017), show that the dynamics of these ice deposits are driven by latitude, elevation, seasonal evolution, and long-term orbital cycles (see a review in Moore, 2021; Young et al., 2021). For example, Bertrand and Forget (2016) and Bertrand et al. (2018) showed that the accumulation of N2 ice occurs within Sputnik Planitia as a result of the lower elevation in the basin, which leads to a stronger thermal infrared cooling that is balanced by a higher condensation temperature of N2. This in turn leads to a higher condensation rate than in any other location. Bertrand et al. (2019) showed that the orbital configurations of Pluto in the past Earth million years would have led to a net accumulation of CH4-rich deposits in the equatorial regions of Pluto, and to seasonal accumulation elsewhere. The persistent absence of perennial ice in the equatorial region of Cthulhu may stem from its albedo, which is exceptionally low compared to the rest of the Pluto surface. This enhances solar heating of the high-inertia water-ice substrate. This stored heat is then released during colder seasons and keeps surface temperatures higher than elsewhere, which in turn inhibits CH4 frost accumulation (Earle et al., 2018a). In contrast, the high albedo of the seasonal CH4 deposits at medium and high latitudes favors surface cooling in winter, promoting N2 condensation and contributing to the formation of N2-rich terrains at these latitudes (Earle et al., 2018a; Bertrand et al., 2019). While these models successfully capture the first-order mechanisms controlling the large-scale ice distribution observed by New Horizons, it remains challenging to accurately quantify the amount of and sublimation rate or condensation because key surface and thermal parameters (e.g., the CH4 ice albedo or the water-ice thermal inertia) are only loosely constrained by the available observations (Young et al., 2021).
@@
Recent analyses using high-resolution measurements by Ralph/MVIC and the Long-Range Reconnaissance Imager (LORRI) instruments revealed bright N2 and CH4 ice-rich deposits on crater floors, rims, and mountain tops (Bertrand et al., 2020; Earle et al., 2022). More specifically, Bertrand et al. (2020) highlighted the bright CH4 frost on the north-facing crater rims and slopes in the Cthulhu region and at the top of Pigafetta and Elcano Montes. Using high-resolution modeling, the authors showed that the methane deposits found at the top of these mountains form from a circulation-induced enrichment of gaseous methane a few kilometers above the Pluto plains that favors methane condensation at mountain summits. Their model cannot be used to study the formation of CH4 deposits on crater rims, however, because the resolution of these craters is too thin compared to the model resolution. The authors suggested that these deposits found on the crater flanks might be linked to the specific microclimates of these slopes. Because the atmosphere of Pluto is tenuous, sloped terrains receive significantly different amounts of solar insolation compared to flat surfaces during winter and spring (see their Figs. S8 and S9). This creates localized cold microclimates on the slope that can favor the accumulation of volatiles and mitigate their sublimation. They therefore suggested that the frosts observed on the Cthulhu crater rims are seasonal frost that formed during the previous winter and now sublimates. Earle et al. (2022) extended this work to other regions of Pluto and found a latitudinal dependence in the orientation of CH4 frost on crater rims. They showed that methane frost is detected on north-facing slopes in the equatorial regions of Pluto ($\pm$ 15° latitude), on equatorward-facing slopes at higher latitudes (between 15° N and 38°N), and then on all of the crater rims at higher latitudes. Although they shared the hypothesis by Bertrand et al. (2020) to explain the methane frost at low latitudes on north-facing slopes, they noted that the change in preferential orientation in regions around 18°N might be explained by different formation timescales (e.g., seasonal at low latitudes, and perennial at high latitudes). Finally, for N2 frosts, they showed that they are preferentially observed on the crater floor up to 38° latitude and on equatorward-facing slopes at higher latitudes.
@@
No numerical model has been able to address the frost formation hypotheses proposed by Bertrand et al. (2020) and Earle et al. (2022) so far. These frosts and their associated slope microclimates occur on spatial scales of only a few kilometers Bertrand et al. (2020), which is far smaller than the typical resolution of GCMs (e.g., hundreds of kilometers, Bertrand et al., 2019). Local models, such as 1D radiative equilibrium models, are also inadequate because the stability of methane frost depends on its atmospheric mixing ratio, which is itself governed by large-scale meteorology that cannot be captured by these simplified models. This numerical gray zone has prevented further investigation into the dynamics of methane frosts on the Pluto slopes so far.
@@
A similar problem was encountered by the Martian community with the modeling of water frost on tropical slopes. Its thin atmosphere is transparent to visible radiation, and slope-driven microclimates are therefore significant on Mars. These slope microclimates allow the formation of CO2 frost (an analog for N2 on Pluto) and water frost (an analog for CH4 on Pluto) at subtropical latitudes, up to 20–30 degrees equatorward of the seasonal polar caps (e.g., Vincendon et al., 2010a; Lange et al., 2024). The same numerical limitations as described above applied when these low-latitude frosts on crater walls were modeled. Recently, Lange et al. (2023) introduced a subgrid parameterization of slope microclimates in the Mars Planetary Climate Model. The principle of this parameterization is illustrated in Fig. 2. For each grid cell of the global Mars climate model, the terrain is decomposed into a distribution of sloped surfaces (defined by characteristic slope angles) and a flat surface. In each subgrid terrain, the surface energy balance and the condensation/sublimation of volatiles are computed under the assumption that all terrains interact with the same atmospheric column. The near-surface atmosphere then experiences an average effect of these subgrid microclimates. This parameterization successfully reproduced the seasonal dynamics of CO2 and water frosts at low latitudes on Mars.
@@
In this study, we adapted the slope microclimate parameterization developed by Lange et al. (2023) for Pluto to (1) describe the seasonal evolution of methane and nitrogen frost on crater slopes over a Pluto year, with particular attention to the sensitivity to thermophysical and optical properties of the ices, (2) discuss the nature of the frosts observed by New Horizons on crater flanks and slopes , and (3) discuss the potential effect of the slope microclimates on the global climate of Pluto. To this end, we describe the parameterization in Sect. 2. The dynamics of methane and nitrogen frost on sloped terrains and the nature of the frosts observed by New Horizons are presented in Sect. 3. The potential implications for the global climate of the slope microclimates are discussed in Sect. 4.1. We draw our conclusions in Sect. 5.
@@ 2.1 Frost identification and statistics
In order to compare our model with the observations of frost on slopes, we first determine the typical slope angle and azimuth at which frost is observed. While the mapping of N2/CH4 frosts has already been made in previous studies (e.g., Schmitt et al., 2017; Earle et al., 2018b), we focus here on the Cthulhu region as it is an unique location where frost is only observed on slopes and not on the dark plains of the regions (Bertrand et al., 2020), suggesting the crucial role of slope microclimates -and thereby slope angle and azimuth- in the formation of these frosts. As in Bertrand et al. (2020), we first map the location of bright deposits on LORRI observations (Hofgartner et al., 2023). As some of the bright locations might be linked to direct illumination, we then compare the locations of these bright deposits to the CH4 band depth map and phase index from the LEISA instrument (Schmitt et al., 2017; Gabasova et al., 2021) to derive the nature of these bright surface -frost or not, and if so, if it is pure N2, CH4 diluted in N2 ice or a CH4-rich ice phases-. Finally, we determine the slope angle and azimuth at which ice is observed using the Digital Evolution Model (DEM, with a 300 m resolution per pixel) from Schenk et al. (2018). While Hapke photometric analyses can, in principle, provide an independent estimate of surface roughness through the mean slope parameter ($\bar{\theta}$), the restricted phase-angle range of New Horizons data leads to strong parameter degeneracies that prevent us from constraining $\bar{\theta}$ (Protopapa et al., 2020). For this reason, and given that only localized applications of alternative photometric models exist (e.g., Mishra et al., 2025), we use the Schenk et al. (2018) DEM as it the most comprehensive and reliable dataset for slope measurements.
@@ 2.2 The Pluto Volatile Transport Model
For this study, we use the Pluto Volatile Transport Model (VTM) described by Bertrand and Forget (2016, see their Methods section), which has been employed in previous investigations of ice dynamics on Pluto (Bertrand and Forget, 2016; Bertrand et al., 2018, 2019). In short, this model computes the surface radiative budget -assuming that the atmosphere is transparent at all wavelengths-, the exchange of heat with the subsurface by conduction, the exchange of heat with the surface by latent heat, and the exchanges of volatile (N2 and CH4) with the atmosphere. To represent horizontal transport after surface exchanges, the mixing ratios of atmospheric species are forced towards their global mean value, using a timescale of 107 s for CH4, 1 s for N2 estimated from 3D GCM studies (Bertrand and Forget, 2016; Forget et al., 2017).
Simulations are performed with a horizontal grid with a grid-point spacing of 1° by latitude and longitude ($\sim$ 21 km at the equator), and a vertical grid dividing the atmosphere into 27 levels ranging from a few meters close to the surface up to 140 km. The initial distribution of the N2 and CH4 reservoirs is obtained following the 30 million years of volatile ice evolution performed by Bertrand et al. (2018) and Bertrand et al. (2019): an infinite amount of N2 ice is placed within Sputnik Planitia, and an infinite reservoir of CH4 ice is placed in the Bladed Terrain Deposits (located between 15°S and 15°N, 140°W and 15°E).
The emissivity and albedo of such ice deposits are adjustable parameters of the model, and set to have an atmospheric surface pressure and CH4 volume-mixing ratio (also called molar mixing ratio, that is, the number of moles of the CH4 gas divided by the number of moles of air, in percent) consistent with the measurements during the New Horizons fly-by (see Sect. 2.4). The albedo of the frost/ice-free terrains are set to observed values from Hofgartner et al. (2023) and their emissivity to 1. The diurnal thermal inertia of the bare-surface, N2 ($\rm{TI}_{\rm{N}_{2},\rm{d}}$) and CH4 ($\rm{TI}_{\rm{CH}_{4},\rm{d}}$) are set by default to 9 J m-2 K-1 s-1/2 following low-thermal inertia derived by Lellouch et al. (2011) and Bertrand et al. (2025). The seasonal thermal inertia of the bedrock $\rm{TI}_{\rm{bedrock},\rm{s}}$ is another adjustable parameter. The albedo of a surface is changed when ice/frost is present at the surface following the albedo-feedback parameterization of Bertrand et al. (2020).